OpenStudy (anonymous):

What is the value of the discriminant of 2x^2 + 3x – 14 = 0? 121 –29 35 45

OpenStudy (anonymous):

x= -2, 7/2

OpenStudy (anonymous):

\[b^2-4ac=D\] 9-4(-28)= 9+112=121